Sky High Company has two departments X and Y. The following estimates are for the coming...

  1. Sky High Company has two departments X and Y. The following estimates are for the coming year:

X

Y

Direct manufacturing labor-hours

20,000

40.000

Machine-hours

40,000

20,000

Manufacturing overhead

$200,000

$400,000

A single indirect cost rate based on direct manufacturing labor-hours for the entire plant is     ?

  1. $12 per direct labor hour
  2. $15 per direct labor-hour
  3. $21 per direct labor-hour
  4. $10 per direct labor hour
  5. None of the above
Answer #1

correct option is"d"- $10 per direct labor hour

Indirect rate per direct manufacturing labor hours =Total overhead of plant /Total direct manufacturing labor hours

                                       = 600000/60000

                                      = $ 10 per direct labor hour

#

Total overhead of plant = 200000 for X+ 400000 For Y = 600000

Total direct manufacturing labor hours= 20000 +40000 = 60000 labor hours